OpenStudy (anonymous):

Group Theory "Using without proof, the homomorphism theorem, or otherwise, show that U(N)/SU(N) is isomorphic to  U(1)." I really struggle with group theory, so if anyone is willing to take me through this slowly , I would be very grateful!

OpenStudy (primeralph):

My worst nightmare as a child: Group Theory.

OpenStudy (anonymous):

@primeralph I know the exact feeling..I just can't get my head around it at all

OpenStudy (primeralph):

@sarahusher I could send you a pretty good book for this, but I can't help. Can barely remember how to solve this.

OpenStudy (kinggeorge):

I'm unfamiliar with the notation U(N), and SU(N). What are these called/defined as?

OpenStudy (anonymous):

what does '"without proof" mean?

OpenStudy (dan815):

whats the name of the book i wanna learn this stuff

OpenStudy (anonymous):

if i am not mistaken (which i may very well be) an element of \(U(1)\) is a complex number \(e^{i\theta}\)

OpenStudy (primeralph):

@dan815 Hold on, it's somewhere in a hard drive; I had it with Lie Theory for a while, back in High School.

OpenStudy (primeralph):

@sarahusher Any idea what the notations mean?

OpenStudy (anonymous):

What part don't you understand - do you need help with the definitions of U(n) and SU(n)?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

U(N) in defined as the Unitary Matrix such that \[U(N)=\left\{ A \in M _{N}(Complex): A ^{t}=A ^{-1} \right\}\] and SU(N) is defined as the Special Unitary Matrix such that: \[SU(N)=\left\{ A \in U(N): detA=1 \right\}\]

OpenStudy (anonymous):

and yes @primeralph that book would be great!

OpenStudy (kinggeorge):

Basically, you want to show the existence of a homomorphism \(\varphi:U(N)\to U(1)\) such that \(\ker(\varphi)=SU(N)\). I think the homomorphism you want to use is \[\begin{aligned}\varphi:U(N)&\to U(1)\\A&\to\det(A)\end{aligned}\]

OpenStudy (anonymous):

again maybe i am confused but the obvious map would send an element of \(U(n)\) to the determinant, and kernel would be the matrices with determinant 1

OpenStudy (anonymous):

better expressed by @KingGeorge

OpenStudy (anonymous):

Consider the determinant map $$\det:\mathrm{U}(n)\to \mathrm{U}(1)$$ The kernel of which is (by definition) SU(n). Apply the first homomorphism theorem.

OpenStudy (anonymous):

Hi @Jack172 Sorry I am here, I'm in China at the moment, and the internet is very tempremental so can't decide if it'll let me on the website or not haha Basically I'm confused by the 'quotient' aspect of it, [ie U(N)/SU(N)] and what U(1) is? and how you would even start this off...

OpenStudy (kinggeorge):

Don't get confused by the quotient, it's just notation. In this case, you don't even have to work with it at all. Just prove that \(SU(N)\) is the kernel of the determinant homomorphism, and you'll be able to apply the fundamental theorem of homomorphisms to immediately prove what you want.

OpenStudy (kinggeorge):

and think of U(1) as just the complex numbers with magnitude 1.

OpenStudy (anonymous):

in reality \(G/M\) is a set of cosets, but you can think of it as putting everything in the same bucket (equivalence class) if they have the same feature modulo \(M\) since here \(M=SU(N)\) i.e. matrices with determinant 1, the the "elements" of \(G/M\) only differ by the determinant okay that was a rather lousy way to put it, but it is late

OpenStudy (anonymous):

@KingGeorge Sorry what do you mean when you say 'prove SU(N) is the kernel of the determinant homomorphism?" [when I say I don't understand Group Theory I really don't!) sorry, if this is basic stuff I'm asking!

OpenStudy (anonymous):

the kernel is the set of elements that are sent to the identity, and the identity in \(U(1)\) is the number 1

OpenStudy (anonymous):

so the kernel is by definition \(SU(N)\)

OpenStudy (kinggeorge):

Alright. So the kernel of a group homomorphism \(\varphi:G\to H\) is defined as \(\ker(\varphi):=\{g\in G\,|\,\varphi(g)=1\in H\}\). In this case, the identity of \(U(1)\) is equal to \(\left[1\right]\). Or just 1.

OpenStudy (anonymous):

you also have to say that the determinant map is a homomorphism, but that is clear by the mantra "the determinant of the product is the product of the determinants"

OpenStudy (kinggeorge):

So you want to show that \(A\in SU(N)\) if and only if \(\det(A)=1\). This is basically true from the definition of \(SU(N)\). Satellite brings up a good point. You may want to show the the determinant is indeed a homomorphism.

OpenStudy (anonymous):

by which @KingGeorge means the map \(\phi (A)=\det(A)\) is a homo

OpenStudy (anonymous):

When you say showing the determinant is a homomorphism, do you just mean that: det(AB)=det(A)det(B) ?

OpenStudy (anonymous):

yes

OpenStudy (kinggeorge):

Yes. Although you can probably take that on faith.

OpenStudy (anonymous):

in general the idea for all such problems is this if you want to show that \(G/H\cong G'\) find a (usually obvious) map \(\psi\) from \(G\to G'\) where everything in \(H\) gets sent to the identity of \(G'\)

OpenStudy (anonymous):

it should be the first thing that springs to mind more or less

OpenStudy (anonymous):

okay, also if U(1) is the complex numbers with magnitude 1 what is an example of SU(2)... I think the notation isn't helping me, cause I can't visualise it

OpenStudy (anonymous):

my understanding would be 2 by 2 unitary matrices with determinant 1

OpenStudy (anonymous):

@KingGeorge seem right?

OpenStudy (kinggeorge):

seems right to me.

OpenStudy (anonymous):

right okay, I get that

OpenStudy (anonymous):

\(U(1)\) is "one by one" unitary matrices, i.e. complex numbers of magnitude 1

OpenStudy (anonymous):

and SU(3) would be a 3x3 unitary matrix with det=1 and so on...?

OpenStudy (kinggeorge):

correct.

OpenStudy (anonymous):

that is my understanding, yes

OpenStudy (anonymous):

and what is \[Z _{2}\], is that 2x2 matrices with integers?

OpenStudy (anonymous):

no, \(\mathbb{Z_2}=\{0,1\}\)

OpenStudy (kinggeorge):

under addition.

OpenStudy (anonymous):

with the operation addition mod 2

OpenStudy (anonymous):

also known as \(\mathbb{Z}/2\mathbb{Z}\)

OpenStudy (anonymous):

probably the simplest ever version of such a proof would be to show that \[\mathbb{Z}/\mathbb{2Z}\cong \mathbb{Z_2}\]

OpenStudy (anonymous):

Okay thankyou so much @satellite73 and @KingGeorge for your help [and patience] :) I'm going to try and work through a few others, and try and get used to proving these things... I think I need to spend a bit of time looking at examples of what each notation means

OpenStudy (kinggeorge):

You're welcome, and good luck.